Jjsksjsxkdmf:

6) Mandy and 2 friends bought some mechanical pencils at a special sale. They divided some of the pencils equally among themselves and then gave 3 to Mandy's little brother. At that time they had 19 pencils left. - 3 - 19 to find the number of pencils p that they bought at the sale. 2. 48 pencils b. 37 pencils € 66 pencils d. 22 pencils Solve the equation-

TutoriaIs:

\[\frac{ p }{ 3 }-2=17\] So let's try to get this to equal: \[\frac{ p }{ 3 }=19\] Do you know what to do first?

TutoriaIs:

Oh the user went offline, here's the work: \[\frac{ p }{ 3 }-2=17\] \[~~~~~+2~~~+2\] Add two to both sides \[\frac{ p }{ 3 }=19\] Put the 19 over 1. \[\frac{ p }{ 3 }=\frac{ 19 }{1 }\] Cross mutiply: \[3\times19=57\] \[1\times p=1p\] \[\frac{ 57 }{ 1p }\] then p=57.